# Proposal for DESC cohesive Deep Drilling Programs for LSST A Cohesive Deep Drilling Field Strategy for LSST Cosmology

### v1

@@ 5.5 Combining metric results
A quick analysis of the metric results (Figs 2-4) tend to show that the strategies leading to highest SMoM values do not all fulfill photo- and WL requirements. It may be interested to study in more details these strategies both to quantify to which extent the requirements are not met and propose adjustments of the strategies designed in this paper.
@@
Observations used for photo- calibration (reference catalogs listed in subsubsection 2.2.1) overlap with COSMOS, XMM-LSS and CDF-S. Mean values for these three fields are given in Table 8. A lower depth with respect to reference observations would not be optimal for photo- calibration (a fraction of COSMOS2020 galaxy sample of 89 and 74 would be used for -0.2 and -0.5 mag, respectively) but the final impact has to be estimated using more realistic simulations of the strategies proposed in this paper.
@@ Appendix A Impact of the filter load strategy on SNe Ia metrics
(,) randomly distributed using the G10 intrinsic scatter model (Scolnic & Kessler (2016)) to estimate NSN.
@@ Appendix B DESC cohesive DDF strategies
Moon phase 20% | 480 | 28.3 |
Moon phase 20% | 930 | 24.3 |
@@ Appendix C Simulation results
Results of the PZ metrics are given on Table 12 for the first year of the survey. PZ requirements are fulfilled for all the fields/band but EDFS-a andEDFS-b (differences for and bands for COSMOS are due to a one night difference in the simulations w.r.t. the requirements). Differences observed for EDFS-a and EDFS-b are explained by the fact that these fields are not as deep as other DDFs (SCOC phase 2 recommendations).
@@
We have performed a simulation of the WFD survey using the LSST simulation baseline_v3.0_10yrs. We have simulated a sample of SNe Ia corresponding to 10 times the number of expected SNe Ia (we use the rate from Hounsell et al. (2018)) in the full redshift range [0.01,0.7]. We use the G10 intrinsic scatter model (Scolnic & Kessler (2016)) where (,) distributions are described by asymmetric gaussian distributions with three parameters. We choose random values spanning over the season duration of a group of observations.
SNe Ia fulfilling the selection criteria defined in Table 4 make-up the sample of well-measured SNe Ia used for cosmology measurements. About 900,000 SNe Ia are expected to be observed after ten years (Figure 11) with a rate between 70,000 and 120,000 SNe Ia per year. About 10-15 of the observed SNe Ia lead to accurate distance measurements ().
@@ Appendix D Metric results
SMoM values displayed in Figure 4 are estimated using a random distribution () for the redshifts of the SNe Ia observed in the WFD sample. This choice reflects the way TiDES will observe the sky ((Swann et al., 2019)) but it leads to samples with a low number of SNe Ia at low redshifts (). We have processed the SMoM metric on samples maximizing the number of low SNe Ia of the WFD survey. An increase of SMoM is of 12-13 is observed.

### v2

@@ 5.5 Combining metric results
A quick analysis of the metric results (Figs 2-4) tend to show that the strategies leading to highest SMoM values do not all fulfill photo- and WL requirements. It may be interesting to study in more details these strategies both to quantify to which extent the requirements are not met and propose adjustments of the strategies designed in this paper.
@@
Observations used for photo- calibration (reference catalogs listed in §2.2.1) overlap with COSMOS, XMM-LSS and CDF-S. Mean values for these three fields are given in Table 8. A lower depth with respect to reference observations would not be optimal for photo- calibration (a fraction of COSMOS2020 galaxy sample of 89 and 74 would be used for -0.2 and -0.5 mag, respectively) but the final impact has to be estimated using more realistic simulations of the strategies proposed in this paper.
@@ Appendix A Impact of the filter load strategy on SNe Ia metrics
(,) randomly distributed using the G10 intrinsic scatter model (Scolnic & Kessler, 2016) to estimate NSN.
@@ Appendix B DESC cohesive DDF strategies
Moon phase 30% | 480 | 28.3 |
Moon phase 30% | 930 | 24.3 |
@@ Appendix C Simulation results
Results of the PZ metrics are given on Table 12 for the first year of the survey. PZ requirements are fulfilled for all the fields/bands but EDFS-a and EDFS-b (differences for and bands for COSMOS are due to a one night difference in the simulations w.r.t. the requirements). Differences observed for EDFS-a and EDFS-b are explained by the fact that these fields are not as deep as other DDFs (SCOC phase 2 recommendations).
@@
We have performed a simulation of the WFD survey using the LSST simulation baseline_v3.0_10yrs. We have simulated a sample of SNe Ia corresponding to 10 times the number of expected SNe Ia (we use the rate from Hounsell et al., 2018) in the full redshift range [0.01,0.7]. We use the G10 intrinsic scatter model (Scolnic & Kessler, 2016) where (,) distributions are described by asymmetric gaussian distributions with three parameters. We choose random values spanning over the season duration of a group of observations.
SNe Ia fulfilling the selection criteria defined in Table 4 make up the sample of well-measured SNe Ia used for cosmology measurements. About 900,000 SNe Ia are expected to be observed after ten years (Figure 11) with a rate between 70,000 and 120,000 SNe Ia per year. About 10-15 of the observed SNe Ia lead to accurate distance measurements ().
@@ Appendix D Metric results
SMoM values displayed in Figure 4 are estimated using a random distribution () for the redshifts of the SNe Ia observed in the WFD sample. This choice reflects the way TiDES will observe the sky (Swann et al., 2019) but it leads to samples with a low number of SNe Ia at low redshifts (). We have processed the SMoM metric on samples maximizing the number of low- SNe Ia of the WFD survey. An increase of SMoM of 12-13 is observed.